The length of a rectangle is 5 times 2 less than the width. The perimeter of the rectangle 100 feet. What is the width?
Answer by mananth(16946)

The length of a rectangle is 5 times 2 less than the width. The perimeter of the rectangle 100 feet. What is the width?
...
Let width be w
Length = 5(w-2)
2(L+W)= perimeter
2*(w+5(w-2))=100
2*(w+5w-10)=100
/2
6w-10=50
6w=60
w=10
w=10feet
Length = 5*(10-2)
Length = 5*8
Length = 40 feet
...
m.ananth@hotmail.ca